ANALYSIS OF INTENSITIES OF SPECTRAL FEATURES OF GASEOUS NO IN VIBRATION-ROTATIONAL BAND 6-0 IN ABSORPTION

Y. P. LEE and S. L. CHEAH, Department of Chemistry, National Tsing Hua University, Hsinchu 30013, Taiwan; \underline{J. F. OGILVIE}, Escuela de Quimica, Universidad de Costa Rica, San Pedro, San Jose, Costa Rica.